OpenStudy (loser66):

Show that the ring M of n x n matrices over Real has no ideals other than 0 and M Please, help

OpenStudy (loser66):

I have a guidance but don't get. Let me post

OpenStudy (helder_edwin):

Let's suppose that \(I\neq0\) is an ideal other than \(M\).

OpenStudy (loser66):

Let do the smallest case first as guidance said to easily follow, please

OpenStudy (helder_edwin):

Now let's consider the basis \(\{E_{i,j}\}\) as given in the hint ---this is the cannonical basis of \(\mathbb{R}^{n\times n}\). Which means that given any matrix A we can find \(\alpha_{i,j}\) such that \[ A=\sum_{i,j}\alpha_{i,j}E_{i,j}. \] Actually \(\alpha_{i,j}\) are the entries of A. All right this far?

OpenStudy (loser66):

yes

OpenStudy (misty1212):

the most prosaic way to think about this is if your matrix has one non zero element, then you can multiply it by something to put a non zero element in any entry since the ideal swallows up the product under multiplication, the idea must the the entire matrix ring

OpenStudy (loser66):

for example if \(A =\left[\begin{matrix} 1&2\\3&1\end{matrix}\right]\), then \(A= 1\left[\begin{matrix} 1&0\\0&0\end{matrix}\right]=2\left[\begin{matrix} 0&1\\0&0\end{matrix}\right]+3\left[\begin{matrix} 0&0\\1&0\end{matrix}\right]+1\left[\begin{matrix} 0&0\\0&1\end{matrix}\right]\), right?

OpenStudy (misty1212):

the rest is just matrix multiplication notation which is good to struggle through but not fascinating

OpenStudy (helder_edwin):

Now there is a problem: the hint says to prove that \[\large E_{i,j}E_{k,l}=\delta_{j,k}E_{i,l} \] This is not true!!

OpenStudy (loser66):

*+ before 2, not =

OpenStudy (loser66):

That is true, it leads to 0 is ideal

OpenStudy (helder_edwin):

\[\large E_{2,2}\cdot E_{3,2}= \begin{pmatrix} 0 & 0 & 0\\ 0 & 1 & 0\\ 0 & 0 & 0 \end{pmatrix}\cdot \begin{pmatrix} 0 & 0 & 0\\ 0 & 0 & 0\\ 0 & 1 & 0 \end{pmatrix} =0\]

OpenStudy (loser66):

because the basis \(E_{ij}\) is independent, hence \(E_{ij}E_{kl}=0\)

OpenStudy (loser66):

so that either \(\delta_{ij}=0\) or \(E_{kl}=0\), but \(E_{kl}\neq 0\), hence only \(\delta_{ij}=0\)

OpenStudy (loser66):

and \(\delta_{ij}\) are entry of ideal, and it =0--> ideal =0, right?

OpenStudy (misty1212):

spose \[ M= \left[ {\begin{array}{cc} a & 0 \\ 0 & 0 \\ \end{array} } \right] \] in in the ideal then since \[ \left[ {\begin{array}{cc} a& 0 \\ 0& 0 \\ \end{array} } \right]\times \left[ {\begin{array}{cc} 0& \frac{b}{a} \\ 0 & 0 \\ \end{array} } \right] \] is in the ideal, so is \[ \left[ {\begin{array}{cc} 0 & b \\ 0 & 0 \\ \end{array} } \right] \]

OpenStudy (misty1212):

and therefore so is every damned matrix

OpenStudy (misty1212):

once you got one non zero entry, you got them all

OpenStudy (loser66):

@misty1212 M is a left ideal but not a right one--> hence M is not an ideal

OpenStudy (misty1212):

oh my M is a matrix not an ideal i think i am not really getting the idea across though

OpenStudy (helder_edwin):

If \(I\subset M\) is an ideal (left and right) and if X is a matrix in \(I\) with some non-zero entries (say \(X_{j,k}=\beta\neq0\)) then \[\large E_{j,k}X(1/\beta)E_{j,k}=E_{j,k}\in I \]

OpenStudy (loser66):

then?

OpenStudy (helder_edwin):

And using what misty did. u can put 1's in row j and column k.

OpenStudy (loser66):

After awhile, I confused though. hehehe... what is an ideal? and what is the relationship between left/ right ideal with ideal?

OpenStudy (helder_edwin):

\[\large \begin{pmatrix} a & 0\\ 0 & 0 \end{pmatrix}\cdot\begin{pmatrix} 0 & 1/a\\ 0 & 0 \end{pmatrix}=\begin{pmatrix} 0 & 1 \\ 0 & 0 \end{pmatrix} \] \[\large \begin{pmatrix} 1/a & 0\\ 0 & 0 \end{pmatrix}\cdot\begin{pmatrix} a & 0\\ 0 & 0 \end{pmatrix}=\begin{pmatrix} 1 & 0 \\ 0 & 0 \end{pmatrix} \] \[\large \begin{pmatrix} 0 & 0\\ 1/a & 0 \end{pmatrix}\cdot\begin{pmatrix} a & 0\\ 0 & 0 \end{pmatrix}=\begin{pmatrix} 0 & 0 \\ 1 & 0 \end{pmatrix} \]

OpenStudy (helder_edwin):

all these r in \(I\)

OpenStudy (helder_edwin):

an ideal is both left and right ideal.

OpenStudy (loser66):

and they not need equal, right? just being left and right ideal at the same time,right?

OpenStudy (helder_edwin):

yes

OpenStudy (loser66):

Thank you.

OpenStudy (helder_edwin):

I can't figure out a way to put \(E_{2,2}\) in \(I\), though.

OpenStudy (loser66):

:(

OpenStudy (helder_edwin):

sorry. gotta go.

OpenStudy (loser66):

Thanks for response. I will let my argument here, please check if you have time.